find the value of "a" such that the graphs of 4y = ax+2 and 1/5y = 1/4x+ 7 are parallel
Answer by ilana(305)

Parallel lines have the same slope. If we put the lines in slope intercept form, y=mx+b, then m should be the same in both. 4y = ax+2, so y=(a/4)x+1/2.
1/5y = 1/4x+ 7, so y=(5/4)x+35. So a/4=5/4, so a=5.